a) y^2-11y+28
b) 4x^2-11x+6

Answer by checkley71(8403)

Y^2-11Y+28=0
(Y-7)(7-4)=0
Y-7=0
Y=7 SOLUTION
Y-4=0
Y=4 SOLUTION
------------------------------------------
4X^2-11X+6=0
(4X-3)(X-2)=0
4X-3=0
4X=3
X=4/3 SOLUTION
X-2=0
X=2 SOLUTION